OpenStudy (shrek_wazowski):

-==||WILL FAN AND MEDAL NEED HELP ASAP!||==-Need help on my math test! (5 Questions) 1. What is the solution to the equation? 7x + 6 = –22 A. x = -64/7 B. x = –4 C. x = -16/7 D. x = 4 2. What is the solution to the equation? 4y – 12 = –4 A. y = 11 B. y = 2 C. y = –4 D. y = –13 3. What is the solution to the equation? 3 - z/4 = 8 A. z = –44 B. z = –20 C. z = 20 D. z = 44 4. What is the solution to the equation? 8 + 4s – 2s = 16 A. s = –12 B. s = –4 C. s = 4 D. s = 12

OpenStudy (shrek_wazowski):

5. What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (photon336):

@Shrek_Wazowski try to post one question at a time. it might be easier for us. what's the first thing you need to do? for \(\color{#0cbb34}{\text{Originally Posted by}}\) @Shrek_Wazowski 5. What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12 \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (shrek_wazowski):

Okay :D

OpenStudy (photon336):

so our first step is that we need to get q by itself have an idea of how to do that?

OpenStudy (shrek_wazowski):

No v.v We just got into this Unit and I'm still confused by all of it..

OpenStudy (photon336):

ok cool.. well we need to get all the q's on the same side. whenever we have an equation what we do to one side we do to the other. to get variables on the same side we must do the opposite operation. if we've got addition we subtract from both sides and vise versa make sense?

OpenStudy (shrek_wazowski):

Okay its making more sense now :D

OpenStudy (photon336):

A much more simple example say we have x+3 = 5 I know, it's looks trivial but see if you can apply what I said to this.

OpenStudy (shrek_wazowski):

So 'x' would be 2in that situation

OpenStudy (photon336):

yes but how did you go about doing that?

OpenStudy (shrek_wazowski):

Did th eoppostie of what the equation said so if it was x + 3 = 5 I did x - 3 = 5

OpenStudy (photon336):

exactly. so this is the logic you need to use in order to to these problems

OpenStudy (shrek_wazowski):

So if its multiplying it would be dividing to get tha answer and if it was subtracting it would be adding..

OpenStudy (photon336):

Look at this x-3 = 5 +3 +3 x = ?

OpenStudy (shrek_wazowski):

x equals 2

OpenStudy (mathmale):

Note how Photon is isolating x on the left side. To do this, he is adding 3 to both sides. Can you finish this?

OpenStudy (mathmale):

5 + 3 = ? (It's not 2.)

OpenStudy (shrek_wazowski):

5 + 3 = 8

OpenStudy (photon336):

yeah so that is the point of this @Shrek_Wazowski got x by itself

OpenStudy (shrek_wazowski):

Okay its making a lot more sense now :D

OpenStudy (photon336):

so in the problem you posted \(\color{#0cbb34}{\text{Originally Posted by}}\) @Shrek_Wazowski This is a little more complicated, but we must get all of the Q terms on the same side of the equation, any ideas? 5. What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12 \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (shrek_wazowski):

-4 - 8 = -2 + 16 ?

OpenStudy (photon336):

okay yeah you could subtract 8 from both sides but make sure you do it like this make sure you keep the variables in there too. -4q+8 = -2q-16 -8 +8

OpenStudy (photon336):

sorry that should be -4q+8 - 2q-16-8 -8

OpenStudy (shrek_wazowski):

Okay so you would do -4 + 8 -8 - 2 - 16 - 8

OpenStudy (photon336):

yes! @Shrek_Wazowski but make sure you keep the variable signs in there too so what does this simplify to so far? -4q+8-8 = -2q-16-8

OpenStudy (shrek_wazowski):

Okay so I got -30.. e.o

OpenStudy (photon336):

so 8-8 =0 so that's gone. so we're left with -4q = -2q-16-8 do you know how to subtract negative numbers?

OpenStudy (shrek_wazowski):

Yes so -4q = -2 - 16 - 8?

OpenStudy (shrek_wazowski):

*-2q

OpenStudy (photon336):

YEP keep going

OpenStudy (shrek_wazowski):

Okay so -4q = -26q ?

OpenStudy (photon336):

@Shrek_Wazowski remember you need to incude the variables

OpenStudy (photon336):

-4q = -2q-16-8 -4q = -2q-24

OpenStudy (shrek_wazowski):

so -4 = -2q - 16

OpenStudy (photon336):

-16-8 = -24

OpenStudy (photon336):

so it becomes \(\color{#0cbb34}{\text{Originally Posted by}}\) @Photon336 -4q = -2q-16-8 -4q = -2q-24 \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (shrek_wazowski):

Ooh so -4q = -2q-24

OpenStudy (photon336):

yes :) do you understand it though?

OpenStudy (shrek_wazowski):

80/100 :D

OpenStudy (photon336):

okay so let's bring all the q terms to the same side alright? -4q = -2q-24 we've got -2q right see that -2q we've got to do the opposite we add it to both sides. -4q+2q = -2q+2q-24 -2q = -24

OpenStudy (shrek_wazowski):

Ooh Okay so 2q cancelled out -2q and 2q made -4q into -2q

OpenStudy (photon336):

Yep

OpenStudy (photon336):

how did you get 80/100?

OpenStudy (shrek_wazowski):

Makes a lot more sense now :D and I was saying 80/100 is my level of understanding but I get it now xD

OpenStudy (photon336):

ok :) good so now -2q = -24 so what's q?

OpenStudy (shrek_wazowski):

q = -24?

OpenStudy (photon336):

\[-24 = -2*q \] \[\frac{ -24 }{ -2 } = q \]

OpenStudy (shrek_wazowski):

q = 12

OpenStudy (photon336):

okay so let's do another to solidify this

OpenStudy (shrek_wazowski):

Okay :D

OpenStudy (photon336):

4y-12 = -4 this is from your problems

OpenStudy (shrek_wazowski):

So 4y - 12 = -4

OpenStudy (photon336):

yes that's from your problem set you posted so tell us what do do now and show your work

OpenStudy (shrek_wazowski):

4y - 12 = -4 4y - 12 = 8y to change 8y into -4 you have to subtract So y = -13?

OpenStudy (photon336):

alright so first thing you do is you see - 12 right?

OpenStudy (shrek_wazowski):

Yes

OpenStudy (photon336):

so immediately you should think I need to add +12 to both sides which is the opposite.

OpenStudy (photon336):

so you end up with 4y = -4+12

OpenStudy (shrek_wazowski):

Okay -4 + 12 = 8

OpenStudy (photon336):

4y = 8

OpenStudy (shrek_wazowski):

Ohh so you have to get 4 away from y to get the answer next right?

OpenStudy (photon336):

so you add -4+12 and you get 8 and then you divide both sides by 4 and you get y = 2

OpenStudy (photon336):

Try this one using the strategies we talked about make sure you remember to write in the variables. 7x+6 = -22

OpenStudy (shrek_wazowski):

Okay 7x + 6 = -22 7x + 6 = 13 so thats 13x 13x = -22

OpenStudy (shrek_wazowski):

then you have to get 13 out of the equation to get your answer

OpenStudy (photon336):

wait you cant add 7x+6 because they aren't like terms

OpenStudy (shrek_wazowski):

Oh..

OpenStudy (photon336):

yep they both have to have x in them. you could add them if it were this 7x+6x but you cant add 7x+6 together

OpenStudy (photon336):

@Shrek_Wazowski try again 7x+6 = -22

OpenStudy (shrek_wazowski):

so you have to get x away from seven so you can combine 7 and 6?

OpenStudy (photon336):

you have to get 6 away from 7 :)

OpenStudy (shrek_wazowski):

Okay :p

OpenStudy (shrek_wazowski):

Yes

OpenStudy (photon336):

so you would subtract 6 on both sides and you would be left with 7x = -22-6

OpenStudy (photon336):

7x = -28 so now what would we do

OpenStudy (shrek_wazowski):

Sorry I lost connection..

OpenStudy (photon336):

np so 7x = -28 so now what would we do

OpenStudy (shrek_wazowski):

You would have to get x

OpenStudy (photon336):

\[\frac{ 7x }{ 7 } = \frac{ -28 }{ 7 } = ? \]

OpenStudy (photon336):

I have to go but i'll be on tomorrow if you still need help tag me.